The coordinates below represent two linear equations.

How many solutions does this system of equations have?

Line 1
x y
–4 3
–2 4
Line 2
x y
0 8
2 6


A.
0

B.
exactly 1

C.
exactly 2

D.
infinitely many

Answer by stanbon(75887)

The coordinates below represent two linear equations.
How many solutions does this system of equations have?
Line 1
x y
–4 3
–2 4
slope = (4-3)/(-2--4) = 1/2
-----------------------------------
Line 2
x y
0 8
2 6
slope = (8-6)/(0-2) = 2/-2 = 1
----
Ans: The lines are not parallel so they meet at one point.
One solution
-----------------------
Cheers,
Stan H.
-------------